OpenStudy (lacris):

hi all! Please help me with this math problem! I will give you a shiny medal :) and please give detailed instructions, thanks. If 2/3 is 1/2 of 4/5 of a certain number, then that number is ??

OpenStudy (marcoreus11):

5/3

OpenStudy (marcoreus11):

1/2(4/5)x = 2/3 4/10x = 2/3 x = 2/3 x 10/4 x = 20/12 x = 5/3

OpenStudy (lacris):

ooh I see, so you multiplied 1/2(4/5)x you put the x there to be the unknown number right? and made it equal to 2/3? then divided 2/3 with 4/10 to get x alone then?

OpenStudy (marcoreus11):

yes

OpenStudy (lacris):

oh my gosh, thank you so much! I see now, I was working backwards and didn't think we had to make it equal to 2/3

OpenStudy (marcoreus11):

yup no problem gl

OpenStudy (lacris):

thank you :D

OpenStudy (yanasidlinskiy):

You're welcome